# Towards Goal-oriented Prompt Engineering for Large Language Models: A Survey

### v2

@@ 2.5 Further Selection of Valuable Sub-goals
In the first stage where we decompose a goal into a sub-goal sequence, most existing works rely on LLM’s knowledge to finish the process. Due to the hallucination problem, it is difficult to guarantee that all of the decomposed sub-goals are correct and relevant. To address this problem, we could ask LLMs to try several sub-goals in each step or even several sub-goal sequences, which facilitates the exploration of goal decompositions. Then, we only select those valuable sub-goals from candidates to pursue a global optimal goal achievement. In this section, we introduce methods for such sub-goal postprocessing, which can be divided into three stages of development, Chain-of-thought prompting (CoT) and its variants, Tree of thoughts (ToT) Yao et al. (2023) and its variants, and finally Graph of thoughts (GoT) Besta et al. (2023). An illustration of these approaches is shown in Fig. 2 to better distinguish these methods.
@@
We consider CoT as simply selecting all sub-goals since LLMs have access to all of the sub-goals during the generation process, which makes it prone to irrelevant sub-goals or wrong sub-goal results. Four works are proposed to improve the robustness of CoT through sub-goal postprocess. Self-consistency Wang et al. (2022) prompts LLMs via CoT several times and then conducts majority votes based on the similarity of different outputs from each COT sequence to decide valuable sub-goal sequences. The result from the chosen sequence is considered the final result. MCR Yoran et al. (2023) improves Self-consistency by selecting and combining sub-goals from different sequences. After getting several sub-goal sequences, it uses LLMs to extract valuable sub-goals to form a new sub-goal sequence, dubbed meta-reasoning path, and predict the result based on the meta path. Recmind Wang et al. (2023c) and Selection-Inference Creswell et al. (2022) allow multiple sub-goal candidates at each decomposition step and choose the most valuable one based on LLM self-evaluation.
@@
ToT Yao et al. (2023) advances over CoT by enabling the exploration and comparison of sub-goals based on a tree structure. To make global choices, at each sub-goal when traversing the tree, a decision can be made that involves backtracking or looking ahead based on the implementation of traditional searching algorithms like breadth-first search. For the comparison of sub-goals, ToT solely relies on the self-evaluation of LLMs. CoT and its variants can be considered special cases of ToT, where each node only has one child node.
RAP Hao et al. (2023) and GDP-Zero Yu et al. (2023) improve from ToT’s heuristic-based search to Monte Carlo Tree Search (MCTS). Compared with ToT, MCTS has an additional back-propagation step, where the score value that indicates the goal achievement is back-propagated to update the value score of all sub-goals in the sub-goal sequence. This operation strikes a proper balance between exploration and exploitation to find valuable sub-goals efficiently, which makes MCTS outperform heuristic-based search algorithms on complex or less structured tasks.
@@ 3 Discussion
The 36 representative works in our survey cover 11 tasks, including arithmetic reasoning (11 works), commonsense reasoning (5 works), symbolic reasoning (2 works), logical reasoning (3 works), planning in virtual environment (10 works), planning in real environment (1 work), multihop question answering (5 works), open-domain question answering (1 work), code generation (4 works), dialogue systems (3 works), and recommender systems (1 work). We could see that 60% of the works focus on reasoning, 31% of the works are for planning and the remaining 9% of the works are for other tasks.
For reasoning tasks, all the surveyed works involve only a single stage, with goal decomposition, action execution, sub-goal result evaluation, and valuable sub-goal selection holding 23% equally. There is potential for exploring the synergy of stages, which we will discuss in Section 4. For planning tasks, goal decomposition shares 31%, action selection shares 23%, and sub-goal result evaluation shares the rest 46%. No surveyed work for planning tasks involves the action execution stage and valuable sub-goal selection stage. This is because the currently available evaluation benchmarks are based on either simple real environments or carefully designed virtual environments. Actions are either easy or can be assured to be perfectly executed. Nonetheless, as the field develops, we can expect more complicated environments when actions cost differently or cannot be perfectly executed. Under such scenarios, there should be attention paid to action execution and valuable sub-goal selection stage.
@@
Furthermore, We compare the performance of the surveyed methods in Appendix A (Table 2, 3, 4, 5, 6, 7, 8). Compared with vanilla CoT, on average goal decomposition (Plan-and-solve) improves by 6.1%, action selection (PoT) improves by 2.9%, and valuable sub-goal selection (Self-consistency) improves by 22.7% in arithmetic reasoning, as shown in Table 6. Besides, we observe a gradual increase in performance when integrating additional stages into the method. When combined with Self-consistency, Plan-and-solve can be improved by 11.3% and PoT can be improved by 5.2% on average across several arithmetic reasoning datasets, which indicates that stage synergy may bring further improvement.
During data collection, we find it hard to compare between methods due to the variety of LLMs used, the number of in-context examples included, and the evaluation metric. With our survey, we hope to give researchers a better estimate of the current progress of goal-oriented prompt engineering in different tasks and also promote a unified evaluation procedure for each task for fair comparison of future methods.
@@ 4 Challenges & Opportunities
Existing works only decompose the goal into one layer of sub-goals. However, the sub-goal itself can still be complex and multifaceted. When a certain sub-goal is complex, LLMs may not achieve it perfectly, and the error will propagated, leading to the failure of the final goal. To make sure each sub-goal is simple enough to be handled by LLMs, researchers are now exploring hierarchical decomposition, which further decomposes sub-goals into simpler "sub-sub-goals". Li et al. (2023) is a pioneering work that constructs a dataset for the script generation task focusing on hierarchical decomposition. GITM Zhu et al. (2023) employs the idea of hierarchical decomposition by utilizing a sub-goal tree to connect a sub-goal with even simpler ones. Though a significant success rate increase by employing hierarchical decomposition is reported, relying solely on one tree may not be able to handle more complex problems, which leaves room for the introduction of graphs for further improvement.
@@
We propose to consider two perspectives of efficiency measurement, task-agnostic efficiency and task-specific efficiency. Task-agnostic efficiency aims at measuring the efficiency of model targets for any task, for instance, the total number of prompting times before reaching the sub-goal or final goal. On the other hand, task-specific efficiency is designed to measure efficiency that appears only in certain tasks.
As an example, the number of iterations that a certain method prompts LLMs would affect the speed of plan generation in planning tasks. Such task-agnostic efficiency is important in scenarios like autonomous driving path planning when we need a quick response for better user experiences. On the other hand, the output plan needs to fulfill other criteria, such as minimizing traveling time or distance. Here, the time and distance are task-specific efficiency.
@@
Overall there is a lack of systematic synergy of stages in any task. We found two works that combine their methods, on one stage, with a method, on another stage, in their ablation studies and show performance improvement Chen et al. (2022); Wang et al. (2023b). This suggests the power of synergizing the stages in the goal-oriented framework and careful integration of stages would have the potential for further improvement.
We take the arithmetic reasoning task as an example. To solve a math problem, a naive solution would be to follow Plan-and-solve Wang et al. (2023b) to decompose the math problem into a few logical steps. Then, following MWP Zhang et al. (2023), an operation prediction module can be applied to determine the correct mathematical operations. LLMs can select the operands for performing the operation. Then, LLMs could call an external calculator to calculate the equation to ensure an accurate result. With the result, we could follow Reflexion Shinn et al. (2023) to ask LLMs to double-check the intermediate steps again. We could further repeat the described process to get multiple sequences of sub-goals and follow MCR Yoran et al. (2023) to select valuable sub-goals for more accurate answers. Of course, a naive combination of methods from all stages is heavy, requiring additional attention paid to efficiency.
Additionally, it is also important to consider the between-stages interaction. For the same math problem, if we apply a calculator for calculation (action execution) to guarantee a correct answer, we should be careful when designing the result evaluation stage. If we use LLMs for result evaluation, the correct result from the calculator may inversely be considered as wrong due to the hallucination problem.
@@
From our summarization, we observe an absence of study in certain stages for some tasks, such as code generation and recommender system. There is potential to improve current methods by studying the unexplored stages.
In code generation, all of the surveyed methods focus on leveraging sub-goal evaluation to refine the code generation result. Additional refinement may be done by incorporating other stages. Given a description of the desired code snippets, we could first decompose it into several required sub-functions, and then select appropriate algorithms to realize the functionality of the sub-function. As for the implementation stage of these sub-functions, we could search relevant documents of potential well-written function APIs for reference instead of generating the code fully based on LLMs’ knowledge. APIs in external libraries are regularly updated which can perform better than the static knowledge stored in LLMs.
@@
Moreover, goal-oriented prompt engineering has the potential for broader applications, in for example software development and medical diagnosis, beyond the 11 tasks summarized in this paper.
For instance, code review is a crucial but cumbersome task in software development. It ensures the quality and maintainability of code but requires a reviewer to spend a lot of time doing the systematic examination. Automatic code review can help save such significant effort. Due to the coding ability shown by recent LLMs, there is potential to surpass existing static analysis tools in automatic code review if equipped with our goal-oriented framework. We found one primary study Tufano et al. (2024), which focuses on action selection, already shows that reducing the action space by identifying the type of request changes can significantly improve the success rate of ChatGPT. Here we give another possibility which covers more stages of our framework. With reviewing code as the goal, sub-goals can be documentation quality, function reusability, code style consistency, variable naming readability, and so on. Taking variable naming readability as the sub-goal, we can have LLMs select action between checking the self-consistency of the variable names or comparing those names with the pre-defined naming convention (the action selection stage). Subsequently, LLMs can execute the action by either calling an external tool for similarity measurement or following given heuristic rules (the action execution stage). After going through all the sub-goals (i.e. reviewing criteria), LLMs can revise the given code based on reviews from each perspective.
@@ 5 Conclusion
Appendix A Peroformance Comparison
In this section, we compare the performance of surveyed methods under several tasks. Due to the different settings in papers, we cannot directly take the reported metric score. To address this problem, we first collect scores from those published papers. Note that we not only take the score of the proposed methods in those papers but also that of baselines if they appear as one of our surveyed works. Once the settings of the same method in different papers have no overlap with each other, we take the score under the setting that has the most similar performance. For preprint papers, we take their result once their compared baselines are already recorded and the discrepancy is no more than 3%. As a result, the reported performance in our survey is not strictly a fair comparison, and we don’t make conclusions stating which method works best in certain contexts. We mainly aim to show that the introduction of goal-oriented prompt design can consistently bring improvement. The results are shown in Table , 2, 3, 4, 5, 6, 7, 8.
Methods VirtualHome
Zero-shot Planner 46.1
Re-prompting 41.8
SALP 49.0
Table 2: Performance comparison on planning in virtual environment. The evaluation metric is solve rate (%).
Metric Methods CSQA StrategyQA
Solve rate CoT 79.0 75.3
Self-consistency 81.5 81.6
Plan-and-solve 71.9 65.4
F1 MCR - 73.6
Self-consistency - 72.2
Self-ask - 69.3
@@ Appendix A Peroformance Comparison
Table 3: Performance comparison on commonsense reasoning under solve rate (%) and F1 (%).
Methods HumanEval MBPP
CoT 66.5 48.8
Self-refine 65.2 48.8
Reflexion 91.0 77.1
Self-debug - 75.6
INTERVENOR 75.6 69.8
@@
Methods Amazon Review Yelp
CoT 0.7167 0.9794
ToT 0.7059 0.9766
Selection-Inference 0.6892 0.9698
Table 5: Performance comparison on rating prediction in recommender system. The evaluation metric is Mean Absolute Error.
Methods GSM8K SVAMP ASDiv AQuA AddSub MultiArith
CoT 56.4 69.9 74.0 38.9 85.3 83.8
Self-consistency 74.4 86.6 81.9 48.3 93.7 99.3
PoT 57.0 70.8 - 43.9 85.1 92.2
Plan-and-solve 59.3 75.7 - 46.0 92.2 91.8
Methods Symbolic Reasoning Logical Reasoning
last letter concatenation coinflip PrOntoQA
CoT 70.4 99.0 87.8
Self-consistency 73.4 99.5 89.8
Plan-and-solve 75.2 99.6 -
RAP - - 94.2
Table 7: Performance comparison on symbolic reasoning and logical reasoning. The evaluation metric is solve rate (%).
Metric Methods 2WikiMultihopQA MuSiQue HotpotQA
Accuracy CoT 29.8 12.6 -
Least-to-most 29.0 16.8 -
Self-ask 40.1 16.3 -
F1 MCR 67.9 - 57.0
Self-consistency 65.4 - 56.4
Self-ask 63.8 - 50.2
DecomP 64.1 29.3 56.4
Table 8: Performance comparison on multihop question answering under solve rate (%) and F1 (%).

### v3

@@ 3.5 Further Selection of Valuable Sub-goals
The generated sub-goals in stage (1) may be wrong or irrelevant to overall goal achievement. To address this problem, we could ask LLMs to explore several sub-goals in each step or even several sub-goal sequences and then only select those valuable candidates. In this section, we introduce methods for such sub-goal post-processing, which can be divided into three stages of development, CoT and its variants, Tree of thoughts (ToT) Yao et al. (2023b) and its variants, and finally Graph of thoughts (GoT) Besta et al. (2023). An illustration of these approaches is shown in Fig. 2 to better distinguish these methods.
@@
We consider CoT as simply selecting all sub-goals since LLMs have access to all of the sub-goals during the generation process, which makes it prone to irrelevant sub-goals or wrong sub-goal results. Four works are proposed to improve the robustness of CoT. Self-consistency Wang et al. (2022) prompts LLMs several times via CoT and then conducts majority votes based on the similarity of different outputs from each COT sequence to decide valuable sub-goal sequences. The result from the chosen sequence is considered the final result. MCR Yoran et al. (2023) improves Self-consistency by selecting and combining sub-goals from different sequences. After getting several sub-goal sequences, it uses LLMs to extract valuable sub-goals to form a new sub-goal sequence, dubbed meta-reasoning path, and predict the result based on the meta path. Recmind Wang et al. (2023e) and Selection-Inference Creswell et al. (2022) allow multiple sub-goal candidates at each decomposition step and choose the most valuable one based on LLM self-evaluation.
@@
ToT Yao et al. (2023b) advances over CoT by enabling the exploration and comparison of sub-goals based on a tree structure. To make global choices, at each sub-goal when traversing the tree, a decision can be made that involves backtracking or looking ahead based on the implementation of traditional searching algorithms like breadth-first search. For the comparison of sub-goals, ToT solely relies on the self-evaluation of LLMs. CoT and its variants can be considered special cases of ToT, where each node only has one child node.
RAP Hao et al. (2023) and GDP-Zero Yu et al. (2023a) improve from ToT’s heuristic-based search to Monte Carlo Tree Search (MCTS). Compared with ToT, MCTS has an additional back-propagation step, where the score value that indicates the goal achievement is back-propagated to update the value score of all sub-goals in the sub-goal sequence. This operation strikes a proper balance between exploration and exploitation to find valuable sub-goals efficiently, which makes MCTS outperform heuristic-based search algorithms on complex or less structured tasks.
@@ 4 Discussion
The 50 representative works in our survey cover 11 tasks, including arithmetic reasoning (18 works), commonsense reasoning (5 works), symbolic reasoning (3 works), logical reasoning (5 works), planning in virtual/real environment (12 works), multihop question answering (8 works), open-domain question answering (1 work), code generation (4 works), dialogue systems (6 works), and recommender systems (3 works). We could see that 46% of the works focus on reasoning, 24% of the works are for planning and the remaining 30% of the works are for other tasks.
For reasoning tasks, all the surveyed works involve only a single stage, with goal decomposition, action execution, sub-goal result evaluation, and valuable sub-goal selection holding 27% equally. There is potential for exploring the synergy of stages, which we will discuss in Section 5. For planning tasks, goal decomposition and action selection share 27%, and sub-goal result evaluation shares the rest 46%. No surveyed work for planning tasks involves the action execution stage and valuable sub-goal selection stage. This is because the currently available evaluation benchmarks are based on either simple real environments or carefully designed virtual environments. Actions are either easy or can be assured to be perfectly executed. Nonetheless, as the field develops, we can expect more complicated environments where there should be attention paid to action execution and valuable sub-goal selection stage.
@@
Furthermore, We compare the performance of the surveyed methods in Appendix B (Table 2, 3, 4, 5, 6, 7, 8, 9), and there are some noteworthy findings. In arithmetic reasoning, compared with standard I-O prompting, on average goal decomposition (CoT) improves by 22.6%, sub-goal evaluation (Self-refine Madaan et al. (2023)) improves by 21.1%, and valuable sub-goal selection (Self-consistency Wang et al. (2022)) improves by 32.5%. While valuable sub-goal selection improves the most in arithmetic reasoning, the improvement brought by goal decomposition outperforms sub-goal selection in symbolic reasoning, which indicates that the variations of improvement have relations to task characteristics. Besides, we observe a gradual increase in performance when integrating additional stages into the method. PoT Chen et al. (2022) can be considered as a combination of CoT and action execution, it further improves CoT by 14.7% in arithmetic reasoning. Similarly, Sayplan Rana et al. (2023) in planning tasks, equipped with goal decomposition, action selection, and sub-goal evaluation, outperforms solely sub-goal evaluation (LLM-Planner Song et al. (2023)) by 73.3% in excitability. Thus, we could expect further improvement from stage synergy.
During data collection, we find it hard to compare between methods due to the variety of prompt templates, the number of in-context examples included, and the decoding strategies, even if the methods are applied to the same LLM in the same zero-/few-shot setting. With our survey, we hope to promote a unified evaluation procedure for each task for fair comparison of future methods.
@@ 5 Challenges & Opportunities
Existing works only decompose the goal into one layer of sub-goals, ignoring the fact that the sub-goal itself can still be complex and multifaceted. To make sure each sub-goal is simple enough to be handled by LLMs, researchers are now exploring hierarchical decomposition, which further decomposes sub-goals into simpler "sub-sub-goals". Li et al. (2023) is a pioneering work that constructs a dataset for the script generation task focusing on hierarchical decomposition. GITM Zhu et al. (2023) employs the idea of hierarchical decomposition by utilizing a pre-defined sub-goal tree to connect a sub-goal with even simpler ones. Though a significant success rate increase by employing hierarchical decomposition is reported, a pre-defined tree is not always accessible. Besides, relying solely on one tree may not be able to handle more complex problems, which leaves room for the introduction of graphs for further improvement.
@@
We propose to consider two perspectives of efficiency measurement, task-agnostic efficiency and task-specific efficiency. Task-agnostic efficiency aims at measuring the efficiency of model targets for any task, for instance, the total number of prompting times before reaching the sub-goal or final goal. On the other hand, task-specific efficiency is designed to measure efficiency that appears only in certain tasks. As an example, the number of iterations that a certain method prompts LLMs would affect the speed of plan generation in planning tasks. Such task-agnostic efficiency is important in scenarios like autonomous driving path planning when we need a quick response for better user experiences. On the other hand, the output plan needs to fulfill other criteria, such as minimizing traveling time or distance. Here, the time and distance are task-specific efficiency.
@@
Overall there is a lack of systematic synergy of stages in any task. We found two works that combine their methods, on one stage, with a method, on another stage, and show performance improvement Rana et al. (2023); Chen et al. (2022). This suggests the power of synergizing the stages in the goal-oriented framework and careful integration of stages would have the potential for further improvement.
We take the arithmetic reasoning task as an example. To solve a math problem, a naive solution would be to follow Plan-and-solve Wang et al. (2023d) to decompose the math problem into a few logical steps. Then, following MWP Zhang et al. (2023), an operation prediction module can be applied to determine the correct mathematical operations. LLMs can select the operands for performing the operation. Then, LLMs could call an external calculator to calculate the equation to ensure an accurate result. With the result, we could follow Reflexion Shinn et al. (2023) to ask LLMs to double-check the intermediate steps again. We could further repeat the described process to get multiple sequences of sub-goals and follow MCR Yoran et al. (2023) to remove less valuable equations for more accurate answers. Of course, a naive combination of methods from all stages is heavy, requiring additional attention paid to efficiency. Additionally, it is also important to consider the between-stages interaction. For the same math problem, if we apply a calculator for calculation (action execution) to guarantee a correct answer, we should be careful when designing the result evaluation stage. If we use LLMs for result evaluation, the correct result from the calculator may inversely be considered as wrong due to the hallucination problem.
@@
From our summarization, we observe an absence of study in certain stages for some tasks, such as code generation and recommender system. There is potential to improve current methods by studying the unexplored stages. Moreover, goal-oriented prompt engineering has the potential for broader applications beyond the 11 tasks summarized in this paper.
In code generation, all of the surveyed methods focus on leveraging sub-goal evaluation to refine the code generation result. Additional refinement may be done by incorporating other stages. Given a description of the desired code snippets, we could first decompose it into several required sub-functions, and then select appropriate algorithms to implement them. As for the implementation stage of these sub-functions, we could search relevant documents of potential well-written function APIs for reference instead of generating the code fully based on LLMs’ knowledge. APIs in external libraries are regularly updated which can perform better than the static knowledge stored in LLMs.
@@
For tasks beyond the scope of this paper, here we take code review as an example. Code review is a crucial but cumbersome task in software development. It ensures the quality and maintainability of code but requires a reviewer to spend a lot of time doing the systematic examination. Automatic code review can help save such significant effort. Due to the coding ability shown by recent LLMs, there is potential to surpass existing static analysis tools in automatic code review if equipped with our goal-oriented framework. We found one primary study Tufano et al. (2024), which focuses on action selection, already shows that reducing the action space by identifying the type of request changes can significantly improve the success rate of ChatGPT. Here we give another possibility which covers more stages of our framework. With reviewing code as the goal, sub-goals can be documentation quality, function reusability, code style consistency, variable naming readability, and so on. Taking variable naming readability as the sub-goal, we can have LLMs select action between checking the self-consistency of the variable names or comparing those names with the pre-defined naming convention (the action selection stage). Subsequently, LLMs can execute the action by either calling an external tool for similarity measurement or following given heuristic rules (the action execution stage). After going through all the sub-goals (i.e. reviewing criteria), LLMs can revise the given code based on reviews from each perspective.
@@ 6 Conclusion
Appendix A Related Surveys
While several surveys have explored LLM prompting, none have examined it from a goal-oriented perspective or explicitly summarized the task-agnostic internal steps involved in designing and refining prompting procedures. Some surveys summarize studies that focus on specific tasks or problems. Qiao et al. (2022); Huang and Chang (2022); Lu et al. (2023); Yang et al. (2023) focus solely on reasoning tasks and Wang et al. (2024) on planning tasks. Huang et al. (2023a) deals with the hallucination problem of LLMs. Other studies focus on generally applicable prompting but different sub-fields. Amatriain (2024); Sahoo et al. (2024) simply listed out recent techniques without systematic categorization. Yu et al. (2023b); Chu et al. (2023) study CoT and its XoT variants. Liu et al. (2023d) focus on the prompting framework for facilitating the interaction between LLMs with external worlds. Liu et al. (2023c) focuses on methods for learnable continuous prompt, while our work studies discrete prompt. As many of the will-functioned LLMs are not open-sourced (e.g. GPT-3.5, GPT4), discrete prompts became the focus of recent studies on LLM prompting and therefore we choose to focus on discrete prompting.
Appendix B Performance Comparison
In this section, we compare the performance of surveyed methods under several tasks, including arithmetic reasoning (Table 2), commonsense reasoning (Table 3), symbolic reasoning (Table 4), logical reasoning (Table 5), planning (Table 6), multi-hop question answering (Table 7), code generation (Table 8), and recommender systems (Table 9). The empirical results are collected from corresponding papers or other papers where they serve as baselines. Here, the benchmarks we chose are the ones on which the majority of methods are evaluated. For a full list of benchmarks, we maintain it at https://github.com/Alex-HaochenLi/Goal-oriented-Prompt-Engineering. Note that due to variations in experimental settings like prompt template, decoding strategy, and numbers of in-context examples, even the methods that share the same zero-/few-shot setting and LLM may not be fairly comparable. Therefore, these tables only provide a rough trend of performance, and we don’t make conclusions stating which method works best in certain contexts. We mainly aim to show that the introduction of goal-oriented prompt design can consistently bring improvement.
Methods Stage Setting Model GSM8K SVAMP AQuA AddSub MultiArith
Cobbe et al. (2021) Patel et al. (2021) Ling et al. (2017) Hosseini et al. (2014) Roy and Roth (2016)
Zero-shot CoT Kojima et al. (2022) zero-shot text-davinci-003 56.4 69.9 38.9 85.3 83.8
Few-shot CoT Wei et al. (2022) few-shot text-davinci-003 58.4 80.3 48.4 91.6 93.6
Plan-and-solve Wang et al. (2023d) zero-shot text-davinci-003 59.3 75.7 46.0 92.2 91.8
Least-to-most Zhou et al. (2022) few-shot code-davinci-002 38.3 80.3 40.6 - 74.0
PoT Chen et al. (2022) zero-shot text-davinci-003 57.0 70.8 43.9 85.1 92.2
Toolformer Schick et al. (2023) zero-shot GPT-J - 29.4 - - -
PoT Chen et al. (2022) few-shot code-davinci-002 71.6 85.2 54.1 - -
PAL Gao et al. (2023) few-shot code-davinci-002 72.0 79.4 - 92.5 99.2
Faithful CoT Lyu et al. (2023) few-shot code-davinci-002 72.3 83.4 47.2 - 98.8
RAP Hao et al. (2023) few-shot LLaMA-33B 40.0 - - - -
Self-refine Madaan et al. (2023) few-shot text-davinci-003 64.1 67.6 - - -
Self-consistency Wang et al. (2022) few-shot code-davinci-002 78.0 86.8 52.0 91.6 100.0
Metric Methods Stage Setting Model CSQA StrategyQA
Talmor et al. (2018) Geva et al. (2021)
Solve rate Zero-shot CoT Kojima et al. (2022) zero-shot text-davinci-003 65.2 63.8
Plan-and-solve Wang et al. (2023d) zero-shot text-davinci-003 71.9 65.4
Few-shot CoT Wei et al. (2022) few-shot text-davinci-003 78.3 71.2
Few-shot CoT Wei et al. (2022) few-shot code-davinci-002 79.0 73.4
Self-consistency Wang et al. (2022) few-shot code-davinci-002 81.5 79.8
F1 CoT Wei et al. (2022) few-shot code-davinci-002 - 70.0
Self-ask Press et al. (2022a) few-shot code-davinci-002 - 69.3
MCR Yoran et al. (2023) few-shot code-davinci-002 - 73.6
Self-consistency Wang et al. (2022) few-shot code-davinci-002 - 72.2
@@ Appendix B Performance Comparison
Table 3: Performance comparison on commonsense reasoning under solve rate (%) and F1 (%). F1 is computed by treating prediction and ground truth answers as bags of tokens and computing their precision and recall.
Methods Stage Setting Model last letter concatenation coinflip
Wei et al. (2022) Wei et al. (2022)
I-O Prompting - few-shot text-davinci-002 5.8 49.0
Zero-shot CoT Kojima et al. (2022) zero-shot text-davinci-003 64.8 96.8
Plan-and-solve Wang et al. (2023d) zero-shot text-davinci-003 75.2 99.6
Few-shot CoT Wei et al. (2022) few-shot text-davinci-003 70.6 100.0
Few-shot CoT Wei et al. (2022) few-shot code-davinci-002 70.4 99.0
Least-to-most Zhou et al. (2022) few-shot code-davinci-002 94.0 -
Self-consistency Wang et al. (2022) few-shot code-davinci-002 73.4 99.5
Table 4: Performance comparison on symbolic reasoning. The evaluation metric is solve rate (%).
Methods Stage Setting Model FOLIO ProofWriter
Han et al. (2022) Tafjord et al. (2020)
I-O Prompting - zero-shot gpt-3.5-turbo 48.4 36.4
CoT Wei et al. (2022) few-shot gpt-3.5-turbo 54.9 43.6
Logic-LM Pan et al. (2023) few-shot gpt-3.5-turbo 62.7 58.3
LINC Olausson et al. (2023) few-shot gpt-3.5-turbo 62.6 96.4
LINC Olausson et al. (2023) few-shot GPT-4 72.5 98.3
Table 5: Performance comparison on logical reasoning. The evaluation metric is solve rate (%).
Dataset Methods Stage Setting Model Evaluation Metric
Executability Correctness
Home+Office LLM+P Liu et al. (2023a) few-shot GPT-4 0.0 33.3
LLM-Planner Song et al. (2023) few-shot GPT-4 13.3 66.7
Rana et al. (2023) SayPlan Rana et al. (2023) ①②④ few-shot GPT-4 86.6 73.3
VirtualHome Zero-shot Planner Huang et al. (2022b) few-shot GPT2-large 16.4 33.3
Zero-shot Planner Huang et al. (2022b) few-shot Codex 78.4 46.1
Puig et al. (2018) SALP Gramopadhye and Szafir (2023) few-shot GPT2-large 50.8 49.0
Re-prompting Raman et al. (2022) few-shot Codex 98.9 41.8
Table 6: Performance comparison on planning in virtual environment. The evaluation metrics are executability (%) and correctness (%). Executability measures if the generated actions can be correctly parsed in the environment, while Correctness is a human-annotated metric that assesses whether the given steps could accomplish the given task.
Metric Methods Stage Setting Model 2WikiMultihopQA Bamboogle HotpotQA
Ho et al. (2020) Press et al. (2022b) Yang et al. (2018)
Acc. I-O Prompting - few-shot Codex 25.4 17.6 -
CoT Wei et al. (2022) few-shot Codex 29.8 46.4 -
Least-to-most Zhou et al. (2022) few-shot Codex 29.0 - -
Self-ask Press et al. (2022a) few-shot Codex 40.1 60.0 -
F1 CoT Wei et al. (2022) few-shot Codex 67.2 64.7 56.4
Self-ask Press et al. (2022a) few-shot Codex 63.8 64.6 50.2
DecomP Khot et al. (2022) zero-shot Codex 64.1 25.4 49.9
MCR Yoran et al. (2023) few-shot Codex 67.9 66.5 57.0
Self-consistency Wang et al. (2022) few-shot Codex 65.4 65.0 56.4
Table 7: Performance comparison on multihop question answering under accuracy (%) and F1 (%). F1 is computed by treating prediction and ground truth answers as bags of tokens and computing their precision and recall.
Methods Stage Setting Model HumanEval MBPP
Chen et al. (2021) Austin et al. (2021)
Zero-shot I-O Prompting - zero-shot gpt-3.5-turbo 62.2 41.6
Few-shot I-O Prompting - few-shot gpt-3.5-turbo 65.2 40.6
CoT Wei et al. (2022) zero-shot gpt-3.5-turbo 66.5 48.8
Self-refine Madaan et al. (2023) zero-shot gpt-3.5-turbo 65.2 48.8
INTERVENOR Wang et al. (2023a) zero-shot gpt-3.5-turbo 75.6 69.8
Reflexion Shinn et al. (2023) zero-shot GPT-4 91.0 77.1
Self-debug Chen et al. (2023) few-shot code-davinci-002 - 75.6
@@
Methods Stage Setting Model Amazon Review Yelp
Ni et al. (2019) Geng et al. (2022)
I-O Prompting Liu et al. (2023b) - zero-shot gpt-3.5-turbo 1.1897 1.2359
CoT Wei et al. (2022) zero-shot gpt-3.5-turbo 0.8612 1.1673
Selection-Inference Creswell et al. (2022) zero-shot gpt-3.5-turbo 0.7883 1.0009
I-O Prompting Liu et al. (2023b) - few-shot gpt-3.5-turbo 0.7327 1.0016
CoT Wei et al. (2022) few-shot gpt-3.5-turbo 0.7167 0.9794
ToT Yao et al. (2023b) few-shot gpt-3.5-turbo 0.7059 0.9766
Selection-Inference Creswell et al. (2022) few-shot gpt-3.5-turbo 0.6892 0.9698
Table 9: Performance comparison on rating prediction in recommender system. The evaluation metric is Mean Absolute Error. Here I-O prompting means that LLMs are directly prompted for the final prediction.